OpenStudy (anonymous):

How do I do a, division problem that has the answer as a whole number and a mixed numer as a dividen?

OpenStudy (inkyvoyd):

Give an example please. :)

OpenStudy (anonymous):

? / 5/8=48

OpenStudy (inkyvoyd):

x/(5/8)=48 x*(5/8)/(5/8)=48*(5/8) =>multiply both sides by 5/8 x*1=6*5 x=30

OpenStudy (anonymous):

Thank You!